OpenStudy (anonymous):

Using complete sentences, write an indirect proof proving that if x = 40, then 5 over 6x − 3 ≠ 22 help please

OpenStudy (anonymous):

If \[x \neq 40 \] then \[(\exists x \in Z : \frac{ 5 }{ 6x - 3 } = 22) \] The parenthesis is supposed to be curly brace. So that means if x = 1 then \[\frac{ 5 }{ 6(1) - 3 } = 22\] But you get 1.667 which does not equal 22 so you have a contradiction and the proof is done.

OpenStudy (anonymous):

hi what's the answer then?